### v2

# ROSE: Revolutionizing Open-Set Dense Segmentation with Patch-Wise Perceptual Large Multimodal Model

Abstract
Advances in CLIP and large multimodal models (LMMs) have enabled open-vocabulary and free-text segmentation, yet existing models still require predefined category prompts, limiting free-form category self-generation. Most segmentation LMMs also remain confined to sparse predictions, restricting their applicability in open-set environments. In contrast, we propose ROSE, a Revolutionary Open-set dense SEgmentation LMM, which enables dense mask prediction and open-category generation through patch-wise perception. Our method treats each image patch as an independent region of interest candidate, enabling the model to predict both dense and sparse masks simultaneously. Additionally, a newly designed instruction-response paradigm takes full advantage of the generation and generalization capabilities of LMMs, achieving category prediction independent of closed-set constraints or predefined categories. To further enhance mask detail and category precision, we introduce a conversation-based refinement paradigm, integrating the prediction result from previous step with textual prompt for revision. Extensive experiments demonstrate that ROSE achieves competitive performance across various segmentation tasks in a unified framework. Code will be released.
**footnotetext: Work done during internships at NIO.
1 Introduction
Figure 1: Comparison of existing open-set segmentation frameworks. Both (a) and (b) require predefined category inputs, where (a) uses similarity matching to select the target category, while (b) generates object masks according to the given category. Consequently, method (a) can perform dense prediction, while (b) is restricted in referring segmentation. Our approach, however, eliminates the need for predefined category inputs and produces dense predictions directly. ‘emb’: embedding.
Image segmentation is a fundamental task in computer vision, requiring pixel-level understanding and classification of image content. It reflects the fine-grained perceptual capabilities of vision models, which are crucial for accurate object recognition, scene understanding, and autonomous decision-making. Traditional segmentation methods [46, 64, 9, 78, 25] typically rely on fixed, closed training datasets, limiting their ability to recognize novel or unseen objects and restricting their applicability in real-world scenarios. Recent advancements in visual-language models, such as CLIP [59], facilitate the development of open-vocabulary segmentation (OVSeg) methods, which is able to extend the category range of traditional methods. However, these methods still depend on predefined category candidates to determine the objects to segment (Fig. 1a), positioning them more as “selectors” than true “generators.”
Recently, large language models (LLMs), such as LLaMA [65], ChatGPT [51], and GPT-4 [52], demonstrate powerful capabilities for language understanding, reasoning, and interaction  [65, 51, 52, 15, 87], driving the emergence of large multimodal models (LMMs) like LLaVA [45], PaligeMMA [5], and Ferret  [82]. These models integrate visual and linguistic components, offering more flexibility in visual tasks. However, despite their ability to handle flexible inputs, e.g. text descriptions, segmentation LMMs still rely on category prompts to determine segmentation targets (Fig. 1b) and cannot truly “self-generate” free-form categories. The dependence on predefined categories limits the practical applicability of existing CLIP-based open-vocabulary segmentation models and segmentation LMMs for truly open-set scenarios. Thus how to achieve open-set segmentation without requiring predefined category inputs is a major challenge.
Furthermore, another key limitation with current segmentation LMMs is that most of them adopt sparse prediction rather than dense prediction, where only target regions or key objects in the image are segmented. However, dense prediction is also crucial, with broad applications in fields such as medical image analysis and autonomous vehicle perception. It reflects the model’s ability to handle complex object relationships and co-occurrence within images. An intuitive way to adapt a Segmentation LMM (Fig. 1b) for dense prediction is simply stacking <SEG> tokens in the response. However, this naive approach may result in unstable mask generation, particularly when the number of stacked <SEG> tokens increases. As the sequence length increases, stacked <SEG> tokens absorb long-range spatial dependency, which may cause the model to lose focus on local image details that are crucial for accurate segmentation. Therefore, how to achieve stable and efficient dense prediction while preserve fine-grained image details is another crucial challenge.
To address these challenges, we develop a truly “open” Segmentation LMM, termed as ROSE. It eliminates the requirement for predefined category inputs, and directly performs dense segmentation predictions, as illustrated in Fig. 1c. Specifically, to avoid the long-range spatial dependency caused by stacked <SEG> tokens, we propose the Patch-wise Perception Process, which treats each image patch as an independent region of interest (RoI). Through this process, we obtain three components, including objectness score, mask embedding, and category embedding, for subsequent dense mask prediction and open-category generation. The objectness score serves to filter patches, retaining only those with high scores. We then leverage SAM to decode the mask embeddings of the selected patches, which enables our model to achieve dense mask prediction. Additionally, with the category embedding, we developed an instruction-response paradigm that leverages the generative and generalizable capabilities of LLMs to produce open-set category predictions, eliminating dependence on predefined category sets. In this way, the model is able to generate category names in a language-driven way, free from closed-set limitations, thus empowering the model to classify previously unseen objects autonomously.
Moreover, we introduce a conversation-based refinement mechanism to further improve segmentation details and accuracy. This paradigm allows the model to iteratively refine segmentation boundaries and categories based on user-provided text prompts, thereby enhancing the precision of segmentation masks and the accuracy of category identification, especially in complex or ambiguous visual scenes. In extensive experiments, ROSE achieves competitive performance across various segmentation benchmarks, demonstrating its effectiveness and flexibility as an open-set dense segmentation solution. ROSE sets a new direction for future open-set segmentation in diverse and dynamic environments. In summary, our contributions are as follows:
  • We present ROSE, an innovative Segmentation LMM framework that pioneers the use of patch-wise perception, enabling LMM to perform both dense and sparse mask predictions for the first time.
  • We propose an instruction-response paradigm, fully exploiting the generative and generalizable capabilities of LLMs to achieve open-category generation.
  • With extensive experiments, we demonstrate the effectiveness of ROSE across various segmentation tasks. Additionally, a conversation-based refinement mechanism is introduced that can iteratively improve the accuracy of segmentation boundary and category prediction, particularly in complex or ambiguous visual scenes.
2 Related works
Method Input Linguistic Prompt Dense Segmentation
Image Language Region Class Cands. Ref. Desc. Task Desc. Prediction Refinement
[…]
Table 1: Comparison of the capabilities of classical models. Language denotes the model accepts language modal as input. Region represents that the model can handle regional information. Linguistic Prompt indicates which kind of linguistic information is acceptable. In which, Class Cands. means class candidates provided by humans, Ref. Desc. means instance-level referring description, and Task Desc. means task-level description. Dense Prediction denotes the ability to predict all targets of interest at once.
2.1 Generic Segmentation
Semantic Segmentation aims to classify each pixel in an image according to its category. Early work FCN [46] uses Conv2D as the last layer and predicts category probabilities for each pixel. Subsequent studies focused on improving contextual understanding, some [9, 10] introduced novel context modules, while others [21, 31, 71, 90] explored self-attention mechanisms to capture pixel-wise dependencies.
Instance Segmentation aims at individually identifying each instance of a target object, earlier methods adopt segmentation and grouping techniques to get object proposals. This approach led to the development of bottom-up segmentation strategies [57, 2], including graph-based methods [24, 20] and selective research algorithms [66]. Later, object proposals from Fast R-CNN [23] leverages for instance segmentation [55, 56, 16]. SOLO [72, 73] advances instance segmentation by directly predicting object masks in each spatial grid.
Referring Segmentation is dedicated to segmenting a specific instance based on a natural language description. The basic principle is to merge as much linguistic information as possible to the visual feature [48, 18, 29, 30]. Previous methods focus on the various attention mechanisms [34, 18] to better incorporate language and vision. Recently, with the success of transformer-based models on vision and language area, some powerful works[58, 92, 41] come off.
2.2 Vision Language Models
The emergence of LLMs [65, 51, 52, 15] has led to notable developments in vision-language modeling, where models are designed to understand both visual and textual inputs. Foundational works [40, 17, 93, 45] focus on aligning visual features with language representations, although they are limited in their applicability to region-level tasks.
Recently, vision-language models such as Kosmos [53] and All-Seeing [70] achieved region grounding by employing bounding box-based formats, while models like GPT4RoI [86] and Ferret [82] introduced region-based encoders for enhanced understanding of visual regions. Furthermore, LISA [38] introduces the <SEG> token for pixel-level referring segmentation, with subsequent models like PixelLM [63] and GSVA [77] extending this approach to multi-target referring segmentation. GLaMM [61] introduced a hierarchical feature pyramid for regional prompting, while CoRes [4] incorporated a CoT procedure to improve contextual understanding in segmentation. VisionLLM [69] develops a set of prompts to handle various visual segmentation tasks.
2.3 Open-set Image Segmentation
Recent methods[22, 19, 80] are built on the MaskFormer framework [12], they generate class-agnostic masks and compare the similarity with text embeddings from models like CLIP [60] and ALIGN [32] to classify these regions. OpenSeg [22] utilizes image-level supervision and scales the training data, while models like Zegformer [19] and ZSSeg [80] improved precision by cropping and refining sub-images before processing them with CLIP. Based on them, GKC [26] enhances vision-text alignment by generating synonyms. OVSeg [43] trains a CLIP adapter to boost the performance. ODISE [79] introduces a strong text-to-image diffusion model to learn the text feature space. FC-CLIP [83] designs an end-to-end framework that uses a single frozen CLIP as the backbone. MAFT+ [33] proposes a collaborative framework to optimize vision-text representation jointly. Recently, PSALM [89] replaced the CLIP vision encoder and the transformer decoder of MaskFormer with a large language model [42].
Figure 2: The architecture of ROSE. (a) In Patch-wise Perception Processes, the vision encoder first encodes the input image and gets patched features, the feature is then concatenated with text instruction and fed into the Large Language model. Then every patch is analyzed by the patch analyzer, generating a mask embedding, a category embedding, and an objectness score. (b) In Patch-wise Mask and Category Decoding Process, patches are first filtered with objectness scores. Then mask embedding is fed into the SAM decoder as a prompt for the patch-corresponding mask. Category embedding is employed to make corresponding category predictions in a generative way.
3 Method
In this section, we first define the task in Sec. 3.1, and then detail our ROSE framework in Sec. 3.2 - Sec. 3.4. Finally, we outline the training objectives in Sec. 3.5.
3.1 Task Definition
Given an image $\mathbf{x}_{\text{img}}$ and a text instruction $\mathbf{x}_{\text{txt}}$, the goal is to complete the segmentation procedure, generating a segmentation mask $\hat{\mathbf{M}}$ and category $\hat{\mathbf{y}}_{\text{txt}}$. Open-vocabulary methods require $\mathbf{x}_{\text{txt}}$ to consist of a set of human-provided candidate categories. These methods predict a list of masks $\hat{\mathbf{M}}\in\mathbb{R}^{N\times H\times W}$ and select $\hat{\mathbf{y}}_{\text{txt}}$ from the candidate set. Existing Segmentation LMMs, on the other hand, typically require $\hat{\mathbf{y}}_{\text{txt}}$ to be an instance-level description, which allows them to predict the corresponding target mask $\hat{\mathbf{M}}\in\mathbb{R}^{H\times W}$ without category information.
Our goal is to propose a Segmentation LMM that can not only segment objects based on human prompts or predefined categories but also autonomously predict dense segmentation masks without any human-provided information. Therefore, we define a new task, Free-vocabulary Segmentation, where the model accepts task-level instructions (e.g., “Can you segment the foreground instance?”), and generates a dense segmentation $\hat{\mathbf{M}}\in\mathbb{R}^{N\times H\times W}$ along with the predicted category $\hat{\mathbf{y}}_{\text{txt}}$ in a generative manner.
3.2 Patch-wise Perception
Embedding-based mask generation, as proposed by LISA [38], offers a solution to mask prediction. However, directly stacking <SEG> tokens proves inadequate for dense object prediction (shown in Tab. 3). Contemporary LMMs typically rely on ViT-based encoders or raw image patches for feature processing, handling image data in a patch-wise manner. Inspired by the SOLO instance segmentation model [72], which divides images into grid-based predictions, our model takes image patches as fundamental prediction units, enabling object detection on a finer scale. Our patch-wise perception process is shown in Fig. 2a. Specifically, each patch predicts the following three components: 1) objectness score, indicates the probability of an object of interest being present within the current patch. 2) Mask embedding, serves as input to the SAM module for mask generation. 3) Category embedding, utilized for subsequent classification tasks. To implement this, we begin by dividing an input image $\mathbf{x}_{\text{img}}$ of size $L$$\times$$L$ into non-overlapping patches of size $p$$\times$$p$, yielding $S$$\times$$S$ patches, where $S=\lfloor\frac{L}{p}\rfloor$. Thus, the maximum number of predictions is $S^{2}$. If a target object’s center falls within the spatial region of a patch located at coordinates $(h,w)$, supervision is assigned to that patch and its surrounding patches. All the $S^{2}$ patches are first passed into vision encoder $\mathcal{F}_{\text{vis}}$, and the resulting visual features, combined with the task instruction $\mathbf{x}_{\text{txt}}$, are further processed by LLM $\mathcal{F}_{\text{llm}}$:
$\hat{\mathbf{y}}_{\text{txt}}=\mathcal{F}_{\text{llm}}(\mathcal{F}_{\text{vis}}(\mathbf{x}_{\text{img}}),\mathbf{x}_{\text{txt}}).\end{aligned}$ (1)
We also proposed a mechanism called super-patch, which clusters patches from nearby, designates specialized detecting roles based on object scale (i.e., small, medium, large) and type (thing vs. stuff categories). This role-based segmentation allows the model to adapt predictions based on object characteristics, enhancing performance across varied visual tasks and object scales.
For patches identified as containing target objects, object embeddings $\mathbf{E}_{\text{obj}}$ are extracted from the last layer of LLM. These embeddings are fed into distinct Multilayer Perceptrons (MLPs) to predict the objectness score, mask embedding $\mathbf{E}_{\text{msk}}$, and category embedding $\mathbf{E}_{\text{cat}}$:
$\text{objectness}=&\;\phi_{\text{obj}}(\mathbf{E}_{\text{obj}}),\\ \mathbf{E}_{\text{msk}}=\phi_{\text{msk}}(\mathbf{E}_{\text{obj}}),&\;\mathbf{E}_{\text{cat}}=\phi_{\text{cat}}(\mathbf{E}_{\text{obj}}).\end{aligned}$ (2)
Beyond the aforementioned three components, we also assign more MLPs to predict the SigLIP [85] embedding, which aims to align with the latent space of the SigLIP [85] text encoder.
Method Semantic Seg Instance Seg Referring Segmentation
ADE-20k COCO RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val(U) test(U)
specialist model
Mask2former [13] 57.7 50.1 - - - - - - - -
[…]
Painter [74] 43.4 - - - - - - - - -
SegGPT [75] 34.4 - - - - - - - - -
Pix2Seq v2 [11] - 38.2 - - - - - - - -
PSALM [89] - - 83.6 84.7 81.6 72.9 75.5 70.1 73.8 74.4
Osprey-7B [84] 29.6 - - - - - - - - -
LISA-7B [38] - - 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6
GLaMM-7B [61] - - 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9
GSVA-7B [77] - - 77.2 78.9 73.5 65.9 69.6 59.8 72.7 73.3
GSVA-13B [77] - - 79.2 81.7 77.1 70.3 73.8 63.6 75.7 77.0
Ferret-7B [82] 31.8 - 87.5 91.4 82.5 80.8 87.4 73.1 83.9 84.8
Ferret-13B [82] - - 89.5 92.4 84.4 82.8 88.1 75.2 85.8 86.3
VisionLLM-7B [69] - 30.6 - - - - - - - -
ROSE-7B 43.2 32.4 78.4 80.9 74.5 70.7 76.1 63.0 73.1 73.6
ROSE-7B + CSR 51.6 34.4 87.0 87.7 85.4 86.8 87.5 85.1 85.3 85.9
Table 2: Comparison with SOTA models on common generic segmentation benchmarks. We evaluate our model on ADE-20k dataset for semantic segmentation, COCO dataset for instance segmentation, and refCOCO/+/g for referring segmentation. denotes the paradigm that the model generates a regional description based on the GT mask, both of the results come from Osprey. denotes the framework that the model predicts masks according to the provided category. The best result and second best result are highlighted in bold and underlined.
3.3 Patch-wise Mask and Category Decoding
After obtaining the objections score, mask embedding $\mathbf{E}_{\text{msk}}$, and category embedding $\mathbf{E}_{\text{cat}}$, we further leverage these components to decode the mask and category. An illustration is shown in Fig. 2b with the detailed process as follows.
Patch Filtering Mechanism. The objectness score serves as a filtering mechanism during inference, allowing the model to prioritize patches with high confidence for further segmentation processing. In order to make $\phi_{\text{obj}}$ converge, we will randomly pick some unsupervised patches as negative samples of objections.
Dense Mask Prediction. Following LISA [38], mask embedding $\mathbf{E}_{\text{msk}}$ is fed into the SAM decoder $\mathcal{F}_{\text{dec}}$ as the text prompt, which generates the final mask prediction:
$\mathbf{f}=\mathcal{F}_{\text{enc}}(\mathbf{x}_{\text{img}}),\hat{\mathbf{M}}=\mathcal{F}_{\text{dec}}(\mathbf{E}_{\text{msk}},\mathbf{f}).\end{aligned}$ (3)
$\mathcal{F}_{\text{enc}}$ is the encoder of SAM, which extracts the image feature $\mathbf{f}$ from input image $\mathbf{x}_{\text{img}}$.
Open-category Generation. In contrast to  [38, 69, 63, 77] concentrate on reasoning and neglect classification and  [67, 89] adopt similarity-comparison paradigm, we employ a generative approach where the model produces category predictions through language generation, independent of predefined category constraints. An intuitive adaptation is employing random or learnable queries inside each Patch Analyzer to generate categories. However, the number of such queries is unpredictable, caused by the uncertainty that both the number of words that make up a category and the number of tokens represent the word (e.g. “staircase” worth three tokens for the tokenizer). This problem could be solved by adding a sufficient number of queries, but this will bring great computing costs.
To address this issue, we treat the category embedding $\mathbf{E}_{\text{cat}}$ as a linguistic feature, allowing us to implement a custom instruction-response paradigm for the classification “USER: <CATEGORY> Please decode this linguistic embedding into a noun concept. ASSISTANT: Sure, it is {category_name}.”
Here, <CATEGORY> is the category embedding $\mathbf{E}_{\text{cat}}$, and {category_name} is supposed to be the specific words of target category as the prediction. Specifically, if there are $N$ activated detecting patches, the shape of the input instruction $\mathbf{I}_{\text{cat}}$ will be $N\times L_{\text{seq}}\times D$, in which $L_{\text{seq}}$ is the length of the input sequence, and $D$ represents the hidden dimension.
$\hat{\mathbf{y}}_{\text{cat}}=\mathcal{F}_{\text{llm}}(\mathbf{E}_{\text{cat}},\mathbf{I}_{\text{cat}}).\end{aligned}$ (4)
In the training stage, $N$ is determined by the number of ground truth (GT), and {category_name} $\mathbf{y}_{\text{cat}}$ is provided in the instruction with an attention mask. In the inference, $N$ is a configurable parameter; the detecting patches with top-$N$ objectness score are kept to produce the final prediction. And {category_name} $\hat{\mathbf{y}}_{\text{cat}}$ is generated in an auto-regressive manner simultaneously for all $N$ targets.
3.4 Conversation-based Segmentation Refinement
Recent chain-of-thought works [37, 76] demonstrate that if more explicit instructions are given, LLMs have the potential to sense the details and correct themselves. If LLM can correct its language wrongness, then LMM might also be able to refine its segment predictions and eventually get better results. Following this thought, we propose the CSR paradigm, in which the model takes in the image $\mathbf{x}_{\text{img}}$, mask $\hat{\mathbf{M}}$, category $\hat{\mathbf{y}}_{\text{cat}}$, and refinement instruction $\mathbf{I}_{\text{ref}}$, using these elements to refine segmentation predictions. If one wants an LMM to refine the result, they must let the LMM understand it. To accomplish this, some work [88] introduced DINO [7] and some [82] use pooling-based processing. However, for the simplicity and maintenance of spatial information, we choose to concatenate images and masks directly.
It is worth noting that mask $\hat{\mathbf{M}}$ and category $\hat{\mathbf{y}}_{\text{cat}}$ are not necessarily used according to the refinement scenarios. Overall, we define three key cases: 1) correct classification with imperfect segmentation. 2) incorrect classification with imperfect segmentation. 3) missed detections. Categories are provided in the second case in instruction, and segment masks are used in the first two situations. A full-zero tensor will be concatenated with the image when the mask is absent. Each case is supported by ten unique instructions, with tailored bounding box information to focus the model’s attention on the target.
3.5 Training Objectives
Our training process optimizes five objective functions: 1) Text generation loss, we use Cross-Entropy loss to supervise $\hat{\mathbf{y}}_{\text{txt}}$ and $\hat{\mathbf{y}}_{\text{cat}}$. As mentioned before, $\mathbf{y}_{\text{cat}}$ is the plain category name, and we design a counting task for $\mathbf{y}_{\text{txt}}$, it looks like “Assistant: There are 5 person, 2 bicycle, 4 car, 2 truck, 1 umbrella in this image.”. 2) Following LISA, our mask prediction loss is composed of Dice loss and Binary Cross-entropy (BCE) loss. With the GT mask $\mathbf{M}$ and category $\mathbf{y}_{\text{cat}}$, these can be formulated as:
[…]
3) For objectness loss $\mathcal{L}_{obj}$ we utilize BCE loss. The GT of the patches with the center of the target around is set to positive, and other patches that are extra-picked are set to negative. 4) SigLIP embedding loss $\mathcal{L}_{sig}$ adopts InfoNCE supervision to learn the latent space of the text encoder from SigLIP. The overall loss is formulated as follows:
[…]
4 Experiment
4.1 Experiment setting
Network Architecture. We use llava-onevision-7b [39] as LMM, which contains SigLIP [85] $\mathcal{F}_{\text{vis}}$ and Qwen [3] $\mathcal{F}_{\text{llm}}$, and adopt ViT-H SAM [36] as $\mathcal{F}_{\text{enc}}$ and $\mathcal{F}_{\text{dec}}$. The projectors $\phi$ used to generate embeddings $\mathbf{E}$ are two-layer MLP with a ReLU [1] activation and hidden dimension of 3584. During training, to preserve the knowledge of llava, we employ LoRA [28] on Qwen $\mathcal{F}_{\text{llm}}$. SigLIP $\mathcal{F}_{\text{vis}}$ and SAM vision encoder $\mathcal{F}_{\text{enc}}$ are both frozen. New added projectors $\phi$ and SAM decoder $\mathcal{F}_{\text{dec}}$ are fully fine-tuned. Additionally, the head layer (lm_head) and token embedding (embed_tokens) of Qwen, and the patch layer (patch_embedding) of SigLIP are also trainable.
Implementation Details. Our implementation is based on deepspeed [62]. 8 NVIDIA A800 GPUs are adopted for training. The optimizer is AdamW [47] with a learning rate of 0.003. We use WarmupDecayLR as the learning rate scheduler, and a linear 1500-iteration warmup is set. The total training iteration is 50k, the per-device batch size is 2, and the gradient accumulation step is 10. The input image is resized to 672 $\times$ 672. There are $48^{2}$ predicting patches, and we select the top 100 patches from it for inference.
Dataset and Task. Our training tasks and datasets are composed of the following: 1) Semantic segmentation, we use ADE20k [91], COCO-Stuff [6], and Mapillary [50] in this task. 2) Instance segmentation, In this task, we only use one dataset, COCO [44], in the training. 3) Referring segmentation, Following LISA, we use RefCLEF, RefCOCO, RefCOCO+ [35], and RefCOCOg [49]. 4) Segmentation Refinement. There are two batches of data: the first follows  [14] randomly added some holes and extra patches, and the second randomly shrinks or stretches the object area and keeps parts of this variation.
Segment Unit Classify Uint ADE-20k COCO
Vanilla stack Along <SEG> 37.2 23.2
Dense stack Along <SEG> 34.4 22.6
Patch-wise Along <REGION> 21.5 4.4
Patch-wise Decode embed 47.5 29.2
Table 3: Ablation study of different segmentation framework.
Patch-design ADE-20k COCO
mAP AP50 AP75
w/o super-patch 47.5 29.2 45.7 31.4
$2$$\times$$2$ 43.0 29.8 45.4 32.3
$3$$\times$$3$ 43.2 32.4 49.2 34.8
Table 4: Ablation study of different super-patch.
4.2 Generic Segmentation
To evaluate the effectiveness of our proposed ROSE, we conduct experiments to demonstrate its capabilities on three common segmentation tasks. The prediction of original ROSE and ROSE with conversation-based segmentation refinement (CSR) is reported in Tab. 2.
Semantic Segmentation. In this task, Painter and SegGPT achieving 43.4 and 34.4 mIoU correspondingly. Osprey and Ferret adopt ground-truth masks to get sentence-based responses and calculate their similarity to the vocabulary list to get the category predictions. They score 29.6 and 31.8 mIoU correspondingly. The direct prediction from ROSE gets the result of 43.2 mIoU, slightly lower than Painter. However, CSR largely boosts performance to 51.6 mIoU and achieves SOTA.
Instance Segmentation. This task is more difficult than semantic segmentation because it requires identity information and confidence score in addition. Pix2Seq v2 and VisionLLM adopt the GT category in the prompt and yield class-agnostic masks, they achieve 38.2 and 30.6 mAP correspondingly. ROSE takes the objectness score from the patch analyzer as confidence score, and has a comparable performance of 34.4 mAP.
Referring Segmentation. This task requires the model to understand linguistic instruction, LMMs that use LMMs have natural advantages on it. Ferret concentrates on regional understanding and performs greatly. Our model shows a comparative referring ability in competition with a larger model (Ferret-13B) and achieves multiple SOTAs in refcoco/+/g datasets.
Conversation-based Segmentation Refinement. This task largely unleashes the potential of LMM and boosts its performance. In the semantic segmentation task, we refine the five worst categories predictions, which bring 8.6 mIoU boosts. In instance segmentation, we pick up 10 under-IoU-threshold predictions in the descending order of objectness score and bring 2.0 mAP improvements. In referring segmentation, CSR brings 12.6 $\pm$ 4.4 gains on average.
Target Modules LoRA Alpha ADE-20k COCO
q_proj,k_proj 16 24.7 19.2
q_proj,k_proj 32 25.1 19.5
all_proj 16 26.3 22.7
all_proj 32 29.1 23.1
Table 5: Ablation study of different LoRA parameters.
Method ADE-20k COCO
mAP AP50 AP75
Mask $\uparrow$ 6.4 $\uparrow$ 3.0 $\uparrow$ 4.0 $\uparrow$ 3.1
Mask + Bbox $\uparrow$ 6.3 $\uparrow$ 3.0 $\uparrow$ 5.2 $\uparrow$ 2.8
Concatenation $\uparrow$ 5.4 $\uparrow$ 3.6 $\uparrow$ 6.4 $\uparrow$ 3.7
Table 6: Ablation study of different refinement paradigms.
4.3 Ablation Study
Segmentation Framework. We explore the effectiveness of different paradigms for dense prediction frameworks, results are shown in the Tab. 3. Vanilla stack means stacking $N$ <SEG> together for $N$ targets. Along <SEG> means generate categories ahead of <SEG>. Our experiment shows it performs badly. Dense stack also performs <SEG> stacking but adds more <SEG> for each GT target. The result shows that simply adding more supervision is not helpful. Patch-wise is used by ROSE, it takes the image patches as the predicting unit. Along <REGION> use <REGION> to indicate each patch and generate the corresponding category, our further investigation finds out it causes mismatches between mask and category. By generating categories and decoding them, ROSE gets a more stable and excellent performance.
Super-patch Design. Using patch-wise prediction and decoding manner classification leads ROSE to good performance. Still, it suffers from bad results on instance segmentation. Inspired by PixelLM [63] we propose the super-patch. With spatial $3$$\times$$3$ patches assembled, we assign 4 patches for small object detection, 3 patches for medium object, and one each for large object and stuff region. Results in Tab. 4 show it drags down the performance on semantic segmentation, it may caused by the insufficient number of stuff detectors. But with such multi-scale role-splitting, instance segmentation gets a great gain. We also conduct a $2$$\times$$2$ super-patch design but the performance gain is limited.
Figure 3: Qualitative results. We show some predictions of ROSE in cross-domain and in-domain scenarios, with generated categories labeled near each target. Please zoom in to see the details. The first row shows the results of images from other domains, including crayon drawings and clip art. The second row shows some predictions of the COCO val set.
Figure 4: Visualization of different refinement mechanisms. The first two columns are ground truth and mask expected to be refined. Concat. denotes concatenate mask with image, and Pred. stands for prediction. Mask and Mask+Box are other methods we try.
LoRA Parameter. We investigate the effect of the target module and LoRA alpha in Tab. 5. It shows as the target modules and LoRA alpha increase, the performance continues to improve. It is worth noting that these models are trained with 20% iterations compared to others.
Refinement Mechanism. To make ROSE understand and refine its past predictions, we conduct several different paradigms. The performance gain is shown in the Tab. 6 Besides the concatenating we mentioned above, following FGVP [81], we also run another two experiments: 1) draw the segment mask on the image, and 2) draw the segment mask and bounding box on the image. However, we find that both of them cause target-shifting problems occasionally (shown in Fig. 4 first two rows), especially when counting on small regions. We can tell such degradation from the COCO dataset, in which there are more small objects. We run the evaluation experiments on a subset with 1k samples for time efficiency.
4.4 Qualitative Results
As depicted in Fig. 3, we present the predictions of ROSE in cross-domain and in-domain scenarios. Experiments demonstrate that the model can autonomously and accurately classify and segment instance objects, both within in-domain (row 2) COCO scenes and in cross-domain (row 1) crayon drawings and clip art scenes.
5 Conclusion
In this paper, we presented ROSE, a novel framework enabling dense mask prediction and open-category generation across the image. We designed the patch-wise perception process, which treats each image patch as an independent region, addressing the dense prediction problem. We also proposed a new instruction-response paradigm, allowing the model to classify in a generative way. To further unleash the power of LMM, we introduced a conversation-based refinement mechanism, which largely boosts the performance. We hope our work shows a creative perspective for the coming open-set segmentation works.
Limitation Our approach advances open-set dense segmentation, but lacking a comprehensive benchmark limits our ability to fully evaluate model performance across diverse open-set scenarios.
Supplementary Material
Appendix A Impelmentation Details
This section introduces more experiment details about ROSE, including dataset and task settings in Sec. A.3, the exact $3$$\times$$3$ super-patch arrangement in Sec. A.2, and the mechanism of refinement script that aims to stimulate human behavior in Sec. A.1.
A.1 Refinement Mechanism
Semantic Segmentation With segmentation result $\mathbf{M}_{sem}$, we first calculate a confusion matrix between GT. Based on the IoU metric, which is the ratio of intersection and union, we propose Union minus Intersection (UmI):
[…]
UmI shows the wrongness of prediction for each category. We then pick the five highest UmI results for further refinement. Finally, the situation and prompt are determined by the recall rate with the following algorithm:
Input: matrix, recall, cat_idx
Output: situation
# matrix (Tensor), confusion matrix, shape (N, N)
# recall (List[Float]), list of recall rate
# cat_idx (List[Int]), list of category to refine
1 for i, idx in enumerate(cat_idx):
2       if recall[i] < 0.2 :
             # matrix[:, n], pixel belong to class n
             # matrix[m, :], pixel predicted as class m
3             other_cat_iou = matrix[:-1, idx].max() / matrix[:, idx].sum()
4            if other_cat_iou > 0.5 :
                   # incorrect classification
5                   situation = “category”
6            else
                   # missed detections
7                   situation = “missed”
8
9      else
             # correct classification
10             situation = “mask”
11
Algorithm 1 Judge Refinement Situation
Instance Segmentation With N instance prediction after processing, we first calculate an IoU matrix with M GT and keep the highest IoU result for each prediction as the matching result. Then, after descending sort by objectness score, we select ten predictions under 50 IoU for the refinement. Finally, the situation is selected from the first two situations, determined by the correctness of the classification result.
Referring Segmentation Because referring segmentation predicts mask solely, we use the first situation (mask) to refine every prediction.
A.2 Super-patch
As mentioned in the main paper, in the default experiment we assign 4 patches for small object detection, 3 patches for medium objects, and one each for large object and stuff region. In Fig. 5 we show how exactly the patches are arranged. A thick gray line indicates the borderline between each super-patch area.
Figure 5: $3$$\times$$3$ super-patch arrangement.
A.3 Dataset and Task
Semantic Segmentation In the training stage, we use instance-level supervision for thing categories, instead of semantic-like supervision. Because we want our model to distinguish different identities as the category may change with the granularity. And instance-level supervision is more reasonable for patch-unit prediction. In the inference, following Mask2former, we stack prediction within the same category and get N-channel mask $\mathbf{M}_{sem}\in\mathbb{R}^{N\times H\times W}$. Here, $N$ is the number of categories of the dataset plus one non-object channel. The prompts we employed look like this “User: <IMAGE> Can you segment this image? Please respond with category names and corresponding segment masks.”.
Instance Segmentation In inference, we first use a threshold to filter predicted instances, and then NMS post-process is conducted to get the final results. And the corresponding prompt like “User: <IMAGE> Please segment all the foreground instances in this image.”.
Referring Segmentation Following LISA, with the {description} annotation in each dataset, one of the prompts we use is shown below “User: <IMAGE> What is {description} in this image? Please output the segmentation mask.”.
Figure 6: Qualitative Results on ADE-20k. Input image, GT, original prediction result, and result after refinement are shown. The corresponding category predictions are tagged on each prediction result.
Segmentation Refinement Here, we show some examples of the prompts used in different situations: 1) the classification is correct but the segmentation mask is corrupted: “User: <IMAGE,MASK> This segmentation mask is incomplete, please ensure the entire object is captured.”. 2) incorrect classification with imperfect segmentation: “User: <IMAGE,MASK> The category of this segmentation result is wrongly predicted as {category}, please correctify this.”. 3) missed detections: “User: <IMAGE,MASK> Please segment target region with mask and corresponding category.”.
Appendix B Qualitative Results
We show more qualitative results here, Fig. 6 shows some results on ADE-20k, and Fig. 7 shows results on cross-domain images and RefCOCO. We pick up various samples to cover all of the pre-defined segmentation refinement situations, which verifies the stability and effectiveness of ROSE and its refinement mechanism.
Figure 7: Qualitative Results on cross-domain images and referring tasks. Prompt, input image, GT, original prediction result and the result after refinement are shown. The corresponding category predictions are tagged on each prediction result for cross-domain samples.